From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Feb 12 2003 - 08:08:39 EST
------------------------------------------------------------ Date: Wed Feb 12 08:08:39 2003, Shift: 00:00-08:00 Supervisor: AIJ, with: IA Log book#14 pages 10 - 11 Shift Summary: ------------------------------------------------------------ 1. Low background for the store period, with one exception when all TPC's and DC's tripped. <br> 2. Achievements : <br> 7868,7869,7870 - Data runs for MRS 90 deg. 1050A and FS 12 deg. 1/8B <br> 7871 - TPC pedestal run <br> 7872 - Non-TPC pedestal run <br> 7873 - DVM run <br> 3. All TPC's and DC's tripped several times (T3 especially). Situation is stable now. ------------------------------------------------------------
This archive was generated by hypermail 2.1.5 : Wed Feb 12 2003 - 08:12:25 EST